What net force is required to cause a 2 kg body to accelerate 4m/s^2?
B.What is the acceleration of an 8 kg mass acted on by a 24 N net force?
C. What is the mass of a body that undergoes an acceleration of 5m/s^2 when acted upon by a 35 N force?

Answer by scott8148(6628)

f = m * a ___ forces are in Newtons , masses are in kilograms , accelerations are in m/s^2

A. f = 2 * 8

B. 24 = 8 * a

C. 35 = m * a